Corso di Matematica - Scienze geologiche - 2003-04 Risultati del primo e secondo compitino

matr. nome  voto # 1 voto # 2 matr. nome  voto # 1 voto # 2
ANONIMO  9  -  Libardi  11  -
Adami  11  -  Limoncelli  17  16
Altomonte  15  -  Loiaconi  7  -
Anfuso  24  11  Maggiore  15  -
Anzilotti  22  12  Mancuso  7  -
Babboni  11  -  Manzi  13  -
Banci  10  11  Marinari  29  22
Beconcini  18  18  Merciadri  10  12
Bendinelli  21  -  Nencini  11  -
Bertoncini  19  -  Nicolai  10  14
Biagioni  15  -  Nicolardi  14  -
Bonaventuri  7  -  Orsetti  27  14
Borzoni  14  20  Orlandini  30  23
Brandani  9  16  Pitanti  10  -
Calò  13  -  Puccetti  19  18
Campinoti  9  -  Puccini  6  -
Camposano  19  14  Renucci  13  15
Ciuffi  9  -  Ridolfi  25  21
Curcio  26  18  Riva  2  -
Dagani  21  12  Rosa  22  19
Damiani  11  6  Rossi  21  6
De Filippo  25  18  Salcioli  19  15
Duranti  7  -  Savio  15  -
Foschi  -  16  Scattareggia  6  -
Garzelli  17  -  Trunfio  17  10
Grossi  3  -  Vagli  15  -
Iuliano  18  16  Ventre  22  -
Landi  20  19